### v1

# Training quantum machine learning model on cloud without uploading the data

Abstract
Based on the linearity of quantum unitary operations, we propose a method that runs the parameterized quantum circuits before encoding the input data. It enables a dataset owner to train machine learning models on quantum cloud computation platforms, without the risk of leaking the information of the data. It is also capable of encoding a huge number of data effectively at a later time using classical computations, thus saving the runtime on quantum computation devices. The trained quantum machine learning model can be run completely on classical computers, so that the dataset owner does not need to have any quantum hardware, nor even quantum simulators. Moreover, the method can mitigate the encoding bottom neck by reducing the required circuit depth from $O(2^{n})$ to $n/2$. These results manifest yet another advantage of quantum and quantum-inspired machine learning models over existing classical neural networks, and broaden the approaches for data security.
I Introduction
Data security is rated more and more important nowadays. Individuals need to keep their privacy from misuse and abuse, companies want to protect their intellectual property rights, and governments concern about its threat to the national security. Unauthorized transfer of data may cause a violation of the law in many places. Some countries also put strict restrictions on exporting data abroad. On the other hand, however, the rapid development of artificial-intelligence technology demands a vast amount of data as input, notably in the training of autonomous driving, intelligent healthcare systems and large language models. At the mean time, not every data owner has very powerful computation devices at home, nor even in his own region. Especially, while quantum machine learning is generally expected to be a potential powerful tool for artificial-intelligence technology, intermediate and large scale quantum computers are available via very few cloud platforms in certain countries only.
To solve the dilemma between the owner of data and the provider of computational resources, here we propose a run-before-encoding method, which can accomplish the following task. Suppose that Alice owns the dataset and Bob holds the quantum cloud computation platform. Alice can run the quantum circuits on Bob’s platform beforehand, without encoding any of her data. Based on the output received from Bob, Alice can input her data later and calculate the cost function needed for training her machine learning model on her own local classical computer. The final trained model can also be run on Alice’s local classical computer. Since her data has never been sent to Bob’s side, it remains perfectly secure against Bob.
As we shall also elaborate below, this method is backed by the linearity in quantum unitary operations. Thus, it is unavailable for existing classical neural networks, where the activation functions of the neurons are nonlinear. Therefore, it displays yet another advantage of using quantum or quantum-inspired machine learning models over existing classical counterparts.
II Typical features of quantum machine learning models
Our method works for quantum circuits with the following features: (1) all the operations in the middle stage are unitary transformations, while the measurements are performed at the last stage only, and (2) the unitary transformations need to result in real probability amplitudes only. Fortunately, the variational quantum circuit (VQC) architecture ml54 ; ml53 ; ml32 widely used for quantum machine learning today belongs to this category, as elaborated below.
Figure 1: A typical variational quantum circuit (VQC) with the RealAmplitudes ansatz.
To avoid confusion, we use $\left|x\right\rangle_{2}$ ($\left|x\right\rangle_{10}$) to denote the quantum states where $x$ is taken as a binary (decimal) number. With this notation, taking the 5-qubit state as an example, we have $\left|0\right\rangle_{10}=\left|0\right\rangle_{2}\otimes\left|0\right\rangle_{2}\otimes\left|0\right\rangle_{2}\otimes\left|0\right\rangle_{2}\otimes\left|0\right\rangle_{2}=\left|00000\right\rangle_{2}$, $\left|1\right\rangle_{10}=\left|00001\right\rangle_{2}$, $\left|2\right\rangle_{10}=\left|00010\right\rangle_{2}$, etc. Fig. 1 shows the typical architecture of a VQC, which consists of three modules: the feature map, the ansatz, and the measurement of the observables. At the beginning, the state of all qubits on the far left are initialized as $\left|0\right\rangle_{10}$. Then the feature map serves as a unitary transformation $U_{F}$, which actually consists of several elemental unitary quantum gates, turning $\left|0\right\rangle_{10}$ into a certain quantum state $\left|x\right\rangle_{10}=U_{F}\left|0\right\rangle_{10}$ that encodes the input data $x$. Currently, there are many data-encoding methods ml118 ; ml269 . But most of them generally have the same feature: the encoded state can be expressed as a linear superposition of the basis vectors. That is, for an input data described by the $d$-dimensional feature vector $x=[x_{0},...,x_{d-1}]^{T}\in R^{d}$, the feature map encodes it using the state
$\left|x\right\rangle_{10}=\frac{1}{C_{x}}\sum\limits_{i=0}^{d-1}f_{i}(x)\left|i\right\rangle_{10},$ (1)
where $f_{i}(x)$ is a function of $x$, $\left|i\right\rangle_{10}$ ($i=0,...,d-1$, $d\leq 2^{n}$) are the basis vectors of the Hilbert space of the $n$-qubit system, and $C_{x}$ is the normalization constant. For example, in the amplitude encoding method, the state has the form
$\left|x\right\rangle_{10}=\frac{1}{C_{x}}\sum\limits_{i=0}^{d-1}x_{i}\left|i\right\rangle_{10},$ (2)
which can be viewed as having taken $f_{i}(x)=x_{i}$. Or in the general qubit encoding method, the state has the form
$\left|x\right\rangle_{10}$ $=$ $\bigotimes\limits_{j=0}^{\left\lceil d/2\right\rceil-1}(f_{j}^{(A)}(x_{2j},x_{2j+1})\left|0\right\rangle_{2}$ (3)
$+f_{j}^{(B)}(x_{2j},x_{2j+1})\left|1\right\rangle_{2}),$
where $f_{i}(x)$ depends not only on a single $x_{i}$.
The ansatz can have many different structure too. One of the widely used one is the RealAmplitudes ansatz ml53 ; ml32 , which is the default ansatz in Qiskit’s VQC implementation. It can have many repetitions and different types of entanglement construction. The example illustrated in Fig. 1 is a 2-repetition one with full entanglement. It starts from a RY gate on each qubits, which introduces an independent rotation angle $\theta$ about the $y$-axis, turning a qubit state $\cos\gamma\left|0\right\rangle_{2}+\sin\gamma\left|1\right\rangle_{2}$ into $\cos\left(\gamma+\frac{\theta}{2}\right)\left|0\right\rangle_{2}+\sin\left(\gamma+\frac{\theta}{2}\right)\left|1\right\rangle_{2}$. Then several CX (controlled-NOT) gates are placed between the qubits to make them entangled, followed by another sets of RY gates on every qubit. This completes one repetition of the RealAmplitudes ansatz. Adding a second repetition means adding another sets of CX gates and RY gates. The effect of the entire RealAmplitudes ansatz can be described by a unitary transformation $U_{A}$. The values of $\theta$ of the RY gates are the adjustable parameters of the ansatz.
A very important merit of the RealAmplitudes ansatz is the following property of the CX gates and RY gates. When being applied on $n$ qubits in the state
$\left|\psi\right\rangle_{10}=\sum\limits_{i=0}^{N-1}A_{i}\left|i\right\rangle_{10},$ (4)
if all $A_{i}$ ($i=0,...,N-1$, $N=2^{n}$) are real numbers, then in the resultant state
$\left|\psi^{\prime}\right\rangle_{10}=\sum\limits_{i=0}^{N-1}A_{i}^{\prime}\left|i\right\rangle_{10},$ (5)
the amplitudes $A_{i}^{\prime}$ ($i=0,...,N-1$) will remain real. That is why the ansatz got its name.
Finally, the measurement module measures a certain observable of the qubits in a certain basis. This can always be rephrased as applying a certain unitary transformation $U_{O}$ on all qubits, then measuring them in the computational basis $\{\left|0\right\rangle_{10},\left|1\right\rangle_{10},\left|2\right\rangle_{10},...,\left|N-1\right\rangle_{10}\}$. This is the only part where non-linearity occurs in a VQC, because unlike unitary transformations, measurements are not linear. The probability distribution obtained from the measurement will then be used for calculating the cost function via classical computations. For example, let $g_{k}$ ($k=1,...,n$) denote the probability for the $k$th qubit to be found as $\left|1\right\rangle_{2}$ in the measurement. Then the widely used mean square error and cross entropy cost functions for each input data $x$ can be calculated as ml94
$L_{MSE}(\lambda)=\frac{1}{n}\sum_{k=1}^{n}(\lambda_{k}-g_{k})^{2}$ (6)
and
$L_{CE}(\lambda)=-\frac{1}{n}\sum_{k=1}^{n}\lambda_{k}\log(g_{k}),$ (7)
respectively, with $\lambda=(\lambda_{1},...,\lambda_{n})$ denoting the label of the input $x$ in the form of one-hot encoding 97ofML94 .
For each input $x$, the final state right before the measurement in the computational basis at the end of the VQC can be expressed as
$\left|\psi_{final}(x)\right\rangle_{10}=\sum\limits_{m=0}^{N-1}a_{m}(x)\left|m\right\rangle_{10},$ (8)
where $a_{m}(x)$ ($m=0,...,N-1$) is the normalized probability amplitude for the basis vector $\left|m\right\rangle_{10}$. If it can be obtained, the probability for finding $\left|m\right\rangle_{10}$ can be computed as
$p_{m}(x)=\left|a_{m}(x)\right|^{2}.$ (9)
In turns, $g_{k}$ can be calculated by summing over any $p_{m}(x)$ where the state of the $k$th qubit in $\left|m\right\rangle_{10}$ is $\left|1\right\rangle_{2}$. Thus, the cost function of an input data $x$ will be achieved.
The rest part is the common procedure for training machine learning models, so we will not go into the details. In brief, by taking the average of the cost functions over all input data, and using the result to further compute the gradients of the adjustable parameters in the ansatz, one epoch of training is completed, just like the training of classical neural networks.
III Our method
Finding $p_{m}(x)$ is the key of the training process. In previous researches, it is done simply by running the VQC by using $x$ as input. If Alice wants to use Bob’s quantum cloud computation platform, she has to send $x$ to Bob. But now we are interested in how to obtain $p_{m}(x)$ without sending $x$ out of Alice’s site. To this end, notice that the relationship between the initial state $\left|0\right\rangle_{10}$, the final state $\left|\psi_{final}(x)\right\rangle_{10}$ before the measurement in the computational basis, and the state $\left|x\right\rangle_{10}$ after applying the feature map is
$\left|\psi_{final}(x)\right\rangle_{10}=U_{O}U_{A}U_{F}\left|0\right\rangle_{10}=U_{O}U_{A}\left|x\right\rangle_{10}.$ (10)
Combining with Eqs. (1) and (8), we yield
$\sum\limits_{m=0}^{N-1}a_{m}(x)\left|m\right\rangle_{10}=\sum\limits_{i=0}^{d-1}\frac{f_{i}(x)}{C_{x}}U_{O}U_{A}\left|i\right\rangle_{10}.$ (11)
Expanding $U_{O}U_{A}\left|i\right\rangle_{10}$ in the computational basis as
$U_{O}U_{A}\left|i\right\rangle_{10}=\sum\limits_{m=0}^{N-1}b_{m}(i)\left|m\right\rangle_{10},$ (12)
then there is
$\sum\limits_{m=0}^{N-1}a_{m}(x)\left|m\right\rangle_{10}=\sum\limits_{m=0}^{N-1}\sum\limits_{i=0}^{d-1}\frac{f_{i}(x)}{C_{x}}b_{m}(i)\left|m\right\rangle_{10},$ (13)
so that
$a_{m}(x)=\sum\limits_{i=0}^{d-1}\frac{f_{i}(x)}{C_{x}}b_{m}(i)$ (14)
for $m=0,...,N-1$. It can be expressed in a more compact form using matrix product as
$\vec{a}(x)=B\vec{f}(x),$ (15)
where $\vec{a}(x)$ is a $N$-dimensional vector defined as
$\vec{a}(x)=[a_{0}(x),...,a_{m}(x),...,a_{N-1}(x)]^{T},$ (16)
$\vec{f}(x)$ is a $d$-dimensional vector defined as
$\vec{f}(x)=\frac{1}{C_{x}}[f_{0}(x),...,f_{i}(x),...,f_{d-1}(x)]^{T},$ (17)
and $B$ is a $N\times d$ matrix defined as
$B=\left[\begin{array}[]{ccccc}b_{0}(0)&\cdots&b_{0}(i)&\cdots&b_{0}(d-1)\\ \vdots&\ddots&\vdots&&\vdots\\ b_{m}(0)&\cdots&b_{m}(i)&\cdots&b_{m}(d-1)\\ \vdots&&\vdots&\ddots&\vdots\\ b_{N-1}(0)&\cdots&b_{N-1}(i)&\cdots&b_{N-1}(d-1)\end{array}\right].$ (18)
Eqs. (12) and (15) give birth to the central idea of our method. The former implies that $b_{m}(i)$ ($m=0,...,N-1$) are the probability amplitudes of the state vector right before the final measurement of the VQC when the basis vector $\left|i\right\rangle_{10}$ ($i=0,...,d-1$) is taken as input. That is, $b_{m}(i)$ is not a function of the data $x$. Consequently, it can be computed without transferring $x$ from Alice to Bob. The latter equation means that Alice can try to obtain all $b_{m}(i)$ ($m=0,...,N-1$, $i=0,...,d-1$) (i.e., the matrix $B$) using Bob’s quantum computation platform first. Then she can decide on the encoding method of her feature map at a later time, which determines the form of the function $f_{i}(x)$. And the normalization constant $C_{x}$ can easily be calculated from $f_{i}(x)$. Finally, she substitutes the values of $b_{m}(i)$ received from Bob into Eq. (15) to calculate $\vec{a}(x)$ and obtain all $p_{m}(x)=\left|a_{m}(x)\right|^{2}$ ($m=0,...,N-1$), which are simple arithmetic that can be done on her own classical computers without even using quantum simulators.
However, the probability amplitude $b_{m}(i)$ is not an observable. Although some quantum simulators can compute it or even find the matrix corresponding to the unitary transformation $U_{O}U_{A}$ directly, it cannot be measured directly on real quantum hardwares. Instead, real quantum computers can output the probability $p_{m}(\left|i\right\rangle_{10})=\left|b_{m}(i)\right|^{2}$ only. Since $b_{m}(i)$ can be a complex number in general, it cannot be uniquely solved from $p_{m}(\left|i\right\rangle_{10})$. That is, it can be expressed as
$b_{m}(i)=s_{m}(i)\sqrt{p_{m}(\left|i\right\rangle_{10})},$ (19)
where $s_{m}(i)$ could be any complex number of unit modulus. Fortunately, as mentioned above, the VQC studied here has the feature that the probability amplitudes $b_{m}(i)$ of the final state before the last measurement are always real. In this case, the possible choices are reduced to either $s_{m}(i)=1$ or $s_{m}(i)=-1$. Still, using $p_{m}(\left|i\right\rangle_{10})$ alone is insufficient for determining $s_{m}(i)$ uniquely.
To circumvent this difficulty, we need to play the following trick. First, by using the basis vectors $\left|i\right\rangle_{10}$ ($i=0,...,d-1$. Note that Eq. (14) indicates that $i=d,...,N-1$ are not needed) as inputs, Alice computes $p_{m}(\left|i\right\rangle_{10})$ ($m=0,...,N-1$, $i=0,...,d-1$) on Bob’s quantum platform. At this stage, from Eqs. (12) and (19) we have
$U_{O}U_{A}\left|i\right\rangle_{10}=\sum\limits_{m=0}^{N-1}s_{m}(i)\sqrt{p_{m}(\left|i\right\rangle_{10})}\left|m\right\rangle_{10}.$ (20)
Next, she picks one of the basis vectors as the reference state $\left|r\right\rangle_{10}$ ($r\in\{0,...,d-1\}$). Then she uses the states $(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})/\sqrt{2}$ ($i=0,...,d-1$, $i\neq r$) as input, and computes $p_{m}(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})$ ($m=0,...,N-1$) on Bob’s quantum platform. Similar to Eq. (20), now we have
$U_{O}U_{A}\left(\frac{1}{\sqrt{2}}(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})\right)$
$=\sum\limits_{m=0}^{N-1}s_{m}(\left|r\right\rangle_{10}$ $+\left|i\right\rangle_{10})\sqrt{p_{m}(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})}\left|m\right\rangle_{10}.$ (21)
Due to the linearity of the unitary transformation $U_{O}U_{A}$, the left-hand side of this equation can be expressed using Eq. (20) as
$U_{O}U_{A}\left(\frac{1}{\sqrt{2}}(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})\right)$ (22)
$=$ $\frac{1}{\sqrt{2}}(U_{O}U_{A}\left|r\right\rangle_{10}+U_{O}U_{A}\left|i\right\rangle_{10})$
$=$ $\sum\limits_{m=0}^{N-1}\frac{1}{\sqrt{2}}(s_{m}(r)\sqrt{p_{m}(\left|r\right\rangle_{10})}$
$+s_{m}(i)\sqrt{p_{m}(\left|i\right\rangle_{10})})\left|m\right\rangle_{10}.$
Comparing with the right-hand side of Eq. (21), we yield
$\frac{1}{\sqrt{2}}\left(s_{m}(r)\sqrt{p_{m}(\left|r\right\rangle_{10})}+s_{m}(i)\sqrt{p_{m}(\left|i\right\rangle_{10})}\right)$ (23)
$=$ $s_{m}(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})\sqrt{p_{m}(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})}.$
Recall that $s_{m}(i)=\pm 1$ which gives $(s_{m}(i))^{2}=1$ and $1/s_{m}(i)=s_{m}(i)$ (so do $s_{m}(r)$ and $s_{m}(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})$), there is
$\left(\frac{1}{\sqrt{2}}\left(s_{m}(r)\sqrt{p_{m}(\left|r\right\rangle_{10})}+s_{m}(i)\sqrt{p_{m}(\left|i\right\rangle_{10})}\right)\right)^{2}$ (24)
$=$ $p_{m}(\left|r\right\rangle_{10}+\left|i\right\rangle_{10}).$
Therefore
$s_{m}(i)=\frac{2p_{m}(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})-(p_{m}(\left|r\right\rangle_{10})+p_{m}(\left|i\right\rangle_{10}))}{2\sqrt{p_{m}(\left|r\right\rangle_{10})}\sqrt{p_{m}(\left|i\right\rangle_{10})}}s_{m}(r).$ (25)
Since $\sqrt{p_{m}(\left|r\right\rangle_{10})}\sqrt{p_{m}(\left|i\right\rangle_{10})}>0$, we have
$s_{m}(i)=\left\{\begin{array}[]{cc}+s_{m}(r),&\left(\begin{array}[]{c}2p_{m}(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})\\ >p_{m}(\left|r\right\rangle_{10})+p_{m}(\left|i\right\rangle_{10})\end{array}\right).\\ -s_{m}(r),&\left(\begin{array}[]{c}2p_{m}(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})\\ <p_{m}(\left|r\right\rangle_{10})+p_{m}(\left|i\right\rangle_{10})\end{array}\right).\end{array}\right.$ (26)
As a result, by comparing $p_{m}(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})$ with $p_{m}(\left|r\right\rangle_{10})$ and $p_{m}(\left|i\right\rangle_{10})$, the relative relationship between the sign of $s_{m}(i)$ and $s_{m}(r)$ can be determined. Consequently, we can write
$s_{m}(i)=\sigma_{m}(i)s_{m}(r)$ (27)
where $\sigma_{m}(i)=s_{m}(r)/s_{m}(i)$ is known to Alice. Combining it with Eqs. (19) and (14), we find
$a_{m}(x)=s_{m}(r)\sum\limits_{i=0}^{d-1}\frac{f_{i}(x)}{C_{x}}\sigma_{m}(i)\sqrt{p_{m}(\left|i\right\rangle_{10})}.$ (28)
Although $s_{m}(r)$ remains unknown, the fact that $(s_{m}(r))^{2}=1$ ensures that we are able to calculate the probability
$p_{m}(x)=\left|a_{m}(x)\right|^{2}=\left(\sum\limits_{i=0}^{d-1}\frac{f_{i}(x)}{C_{x}}\sigma_{m}(i)\sqrt{p_{m}(\left|i\right\rangle_{10})}\right)^{2}.$ (29)
To sum up, by using the state vectors $\left|i\right\rangle_{10}$ ($i=0,...,d-1$) and $(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})/\sqrt{2}$ ($i\neq r$) as inputs (i.e., $d+(d-1)$ states in total), which contain no information on the dataset $x$, Alice can compute the probabilities $p_{m}(\left|i\right\rangle_{10})$ and $p_{m}(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})$ ($m=0,...,N-1$) on Bob’s quantum platform, and calculate $p_{m}(x)$ on her own classical computer later to complete the training of her machine learning model.
Note that if the probability $p_{m}(\left|r\right\rangle_{10})$ for a certain $m$ turns out to be zero when $\left|r\right\rangle_{10}$ is chosen as the reference state, then Eq. (25) will be unable to determine the relative sign of $s_{m}(i)$ ($i\neq r$) because the right-hand side is divided by zero. In this case, Alice needs to pick a second reference state $\left|r^{\prime}\right\rangle_{10}$ ($r^{\prime}\neq r$), and repeat the above procedure by using $(\left|r^{\prime}\right\rangle_{10}+\left|i\right\rangle_{10})/\sqrt{2}$ ($i=0,...,d-1$, $i\neq r^{\prime}$ and $i\neq r$) as input. That is, the VQC needs to be run for an additional $d-2$ times. If the probability $p_{m}(\left|r^{\prime}\right\rangle_{10})$ for the same $m$ is still zero, then Alice has to pick yet another reference state $\left|r^{\prime\prime}\right\rangle_{10}$, until a nonvanishing $p_{m}$ is found. In the worst case, $d-1$ of all the $d$ basis vectors $\left|i\right\rangle_{10}$ will be picked as the reference states and the VQC will be run for $d+(d-1)+(d-2)+...+1=(d+1)d/2$ times. In this case, even if the probabilities $p_{m}$ of these reference states are all zero for the same $m$, the computation can be completed without hurting the result. This is because it means that the value of $a_{m}(x)$ for any $x$ will also be zero for the same $m$, as indicated by Eq. (14). Therefore, the sign of $s_{m}(i)$ ($i=0,...,d-1$) no longer matters for the calculation of $p_{m}(x)$ in this scenario. Luckily, in our experimental test experiment , many sets of the adjustable parameters for the VQC with randomly chosen values are tried and we always picks $\left|r\right\rangle_{10}=\left|0\right\rangle_{10}$ as the only reference state, and it turns out that all the probabilities $p_{m}(\left|r\right\rangle_{10})$ ($m=0,...,N-1$) corresponding to this $\left|r\right\rangle_{10}$ are nonzero. Therefore, running the VQC for $d+(d-1)$ times in total is already sufficient in our experiments.
For clarity, the steps of our method are summarized below.
Protocol 1: The run-before-encoding method
(I) Alice picks the values of the adjustable parameters of the VQC, and sends them to Bob.
(II) Bob sets up the VQC using these parameters, then:
(II-1) He runs the VQC using the basis vectors $\left|i\right\rangle_{10}$ ($i=0,...,d-1$) as input and obtains the output probabilities $p_{m}(\left|i\right\rangle_{10})$ ($m=0,...,N-1$).
(II-2) He picks one of the basis vectors as the reference state $\left|r\right\rangle_{10}$ ($r\in\{0,...,d-1\}$), such that most of the output $p_{m}(\left|r\right\rangle_{10})$ ($m=0,...,N-1$) are nonzero when $\left|r\right\rangle_{10}$ was used as input. Then he runs the VQC using the superposition states $(\left|r\right\rangle_{10}+\left|i\right\rangle_{10})/\sqrt{2}$ ($i=1,...,d-1$, $i\neq r$) as input.
(II-3) If all $p_{m}(\left|r\right\rangle_{10})$ ($m=0,...,N-1$) are nonzero, then the computation on the VQC is completed and Bob goes to step (III). Otherwise, if there is $p_{m}(\left|r\right\rangle_{10})=0$ for at least one specific $m$, he picks a second reference state $\left|r^{\prime}\right\rangle_{10}$ ($r^{\prime}\in\{0,...,d-1\}$, $r^{\prime}\neq r$), such that $p_{m}(\left|r^{\prime}\right\rangle_{10})\neq 0$ for this specific $m$. Then he runs the VQC using $(\left|r^{\prime}\right\rangle_{10}+\left|i\right\rangle_{10})/\sqrt{2}$ ($i=1,...,d-1$, $i\neq r$, $i\neq r^{\prime}$) as input.
(II-4) If there are still some certain $m$ for which both $p_{m}(\left|r\right\rangle_{10})$ and $p_{m}(\left|r^{\prime\prime}\right\rangle_{10})$ are zero, Bob picks a third reference state $\left|r^{\prime\prime}\right\rangle_{10}$ ($r^{\prime\prime}\in\{0,...,d-1\}$, $r^{\prime\prime}\neq r$, $r^{\prime\prime}\neq r^{\prime}$) and runs the VQC using $(\left|r^{\prime\prime}\right\rangle_{10}+\left|i\right\rangle_{10})/\sqrt{2}$ ($i=1,...,d-1$, $i\neq r$, $i\neq r^{\prime}$, $i\neq r^{\prime\prime}$) as input … This procedure is repeated until:
(a) For each $m$, at least one of $p_{m}(\left|r\right\rangle_{10})$, $p_{m}(\left|r^{\prime}\right\rangle_{10})$, $p_{m}(\left|r^{\prime\prime}\right\rangle_{10})$, … is nonzero; or
(b) $d-1$ of all the $d$ basis vectors $\left|i\right\rangle_{10}$ ($i=0,...,d-1$) were picked as the reference states.
(III) Bob calculates the matrix $B$ defined in Eq. (18) using Eq. (19) where $s_{m}(i)$ is determined via Eq. (26). Then he sends $B$ to Alice.
(IV) Alice calculates $\vec{a}(x)$ for every $x$ in her dataset using Eq. (15) on her classical computer. The she computes all $p_{m}(x)=\left|a_{m}(x)\right|^{2}$ ($m=0,...,N-1$), and obtains the cost function as well as the gradients of the adjustable parameters of the VQC. This completes one epoch of training of the quantum machine learning model. For the next epoch of training, Alice adjusts the parameters according to the gradients, and repeats the above procedure from step (I).
IV Advantages
An obvious and foremost advantage of our method is that the security of Alice’s dataset $x$ is well-protected. As can be seen from step (II) of Protocol 1, Bob runs the VQC based on the parameters received from Alice using only the basis vectors and their superpositions as input. None of the data $x$ was encoded in these states, nor being sent to Bob. In fact, the relationship between $x$ and Bob’s output may even remain unknown to Alice herself at this stage, because she does not need to decide on the encoding method and the form of the cost function until step (IV). Therefore, it is clear that Bob does not stand any chance to learn $x$ and the cost function.
Secondly, our method may save the runtime on the quantum devices when the amount of data is huge. As shown in step (II), for example, consider the best case where all the output $p_{m}(\left|r\right\rangle_{10})$ ($m=0,...,N-1$) corresponding to the first reference state $\left|r\right\rangle_{10}$ are nonzero. To obtain the matrix $B$, the VQC needs to be run for $d+(d-1)=2d-1$ input states, regardless the size of the dataset. Once $B$ is obtained, the probability distributions corresponding to all $x$ in the dataset can be calculated via Eq. (15) on classical computers. In contrast, when using the existing conventional method where the VQC needs to be run for each input $x$ one by one, evaluating a dataset with $s$ data points has to call for the VQC $s$ times. Therefore, our method takes a much less occupation time on the quantum devices than the conventional one does for any dataset satisfying $s>>2d-1$. For instance, the MNIST dataset MNIST widely used in machine learning research has a dimension $d=784$, whereas the number of training data is $s=50000$. In this case, our method will be significantly faster. On quantum cloud platforms charging by the runtime, less occupation of the quantum devices also means less cost.
Moreover, our method can help mitigating the encoding bottom neck ml230 . In many encoding methods such as the one described by Eq. (2), encoding a data vector $x$ using $n$ qubits generally requires a circuit depth of the order $O(2^{n})$ ml273 , except for certain sparse vectors. But in our method, the inputs to the VQC are merely basis vectors and simple superpositions of two basis vectors. It can be proven that encoding a single basis vector takes only $n/2$ single-qubit NOT gates in average, while encoding the superpositions of two basis vectors takes only $n/2$ two-qubit controlled-NOT gates in average.
In addition, after the training of the VQC is completed, Alice will no longer need a quantum computer to run it. She can simply use the matrix $B$ obtained in the final epoch of the training as the machine learning model, and applies it on any new input data $x$ by computing Eq. (15) on her local classical computer.
V Further improvements for better security
In Protocol 1, though Bob does not know the value of $x$, he knows the dimension $d$ of $x$, because he is required to run the VQC for $d$ of the basis vectors and their superpositions, while the total number of the basis vectors of a $n$-qubit system is $N=2^{n}$. He also knows the adjustable parameters of the VQC because Alice sent them to him. If these problems are the primary concern while increasing the runtime on the quantum devices is acceptable, we can have the following fix.
To hide the dimension of the data $x$, let us take the MNIST dataset as an example. Originally, Alice needs only to run the quantum circuit for $d=784$ basis vectors instead of all the $N=2^{10}=1024$ basis vectors of $n=10$ qubits. But if Alice wants to blind Bob from knowing that $d=784$, she can ask Bob to run the VQC for all the $1024$ basis vectors instead. Moreover, she can even ask Bob to run a VQC with more than $10$ qubits, which makes it even harder for Bob to guess the actual dimension of her data.
If Alice does not want Bob to know her choice of the adjustable parameters, she can also generate more sets of these parameters whose values are chosen differently by purpose, and ask Bob to run the VQC with these sets. This is like hiding a tree in a forest, so that Bob cannot be sure which set of the parameters is the one that Alice actually uses in her machine learning model.
VI More discussion on the advantage
In conclusion, we propose a run-before-encoding method so that a dataset owner Alice can train a machine learning model on Bob’s quantum cloud computation platform, with the primary advantage that the dataset is kept perfectly secret from Bob. Note that it was pointed out in Ref. ml69 that “almost all branches of quantum machine-learning research have been heavily framed by the question of ‘beating’ classical machine learning in some figure of merit”, but “a number of ‘positive’ results have been put forward that either ‘prove’ theoretically or ‘show’ empirically that quantum computers are better at something”. According to Ref. ml353 , it was even stated at Google I/O 2024 that “no one has yet demonstrated a clear quantum advantage for machine learning on classical data”. But our method works for any quantum circuit where all the operations in the middle are unitary while the measurements are performed at the very end, and the unitary operations always result in real probability amplitudes. These features are not available in existing classical machine learning models but they are not uncommon for quantum ones. Thus, the advantage of our method is neither empirical nor bias towards certain tasks.
Nevertheless, this advantage can be achieved not only on real quantum computers, but also by running quantum simulators on classical computer. That is, though it is an advantage over all existing classical neural networks, rigorously speaking, it may not seem appropriate to regard it as a quantum advantage. We feel that it should be better considered as a “quantum-inspired” advantage. This result also suggests that quantum-inspired classical neural networks, i.e., redesigned classical neural networks by replacing the neuron associated by existing choices of non-linear classical activation functions (e.g., the sigmoid function, ReLU function, and Tanh function, etc.) with unitary linear transformations, should deserve more research interests since it could make full use of the advantage in the near term.
Acknowledgments
This work was supported in part by Guangdong Basic and Applied Basic Research Foundation under Grant No. 2019A1515011048.